OpenStudy (anonymous):

–6p + 7 = 3(2p – 3) – 4(–10 + 4p)

OpenStudy (anonymous):

-6p+7=6p-9+40-16p -6p+7=-10p+31 -6p+10p=31-7 4p=24 p=6